OpenStudy (anonymous):

How many moles of ethane (C2H6) would be needed to react with 62.3 grams of oxygen gas? (4 points) Unbalanced equation: C2H6 + O2 → CO2 + H2O

OpenStudy (anonymous):

@radar @lizzyrockefeller

OpenStudy (anonymous):

I don't want straight answers, I want to learn how to do it, so I can move on & pass my exams in a couple weeks.

OpenStudy (anonymous):

okay il do step by step

OpenStudy (anonymous):

<3 Thanks xD

OpenStudy (anonymous):

your welcome :) it will take me a few okay

OpenStudy (anonymous):

Take your time :) At least your helping ccx

OpenStudy (anonymous):

so what u would do is balance an write the equation like this 2C2H6 + 7O2 --> 4CO2 + 6H2O then you do the molucule weight of O2 which is 32.00

OpenStudy (anonymous):

then after all that you would do 1 Mal of O2=32g 7*32=224g

OpenStudy (anonymous):

im not done im just showing step by step

OpenStudy (anonymous):

whoa... WHOA slow down :3 so first you balance the equation. This means having equal amounts on the product & the reactant. I'm fairly new to this, so I need help with that first. Please?? :3

OpenStudy (anonymous):

okay im sorry

OpenStudy (anonymous):

No worries! Like I said, you're a HUGE help as it is, so anything is appreciated. :3

OpenStudy (anonymous):

yes you have to have equal amounts

OpenStudy (anonymous):

So, to do this, you... multiply? If so, what number would you use, & to what extent?

OpenStudy (anonymous):

you would just do 7*32 to get 224g thats it

OpenStudy (anonymous):

That's it? Are you certain on this? Just want to be sure, not questioning your excellency, just want to get a high grade!

OpenStudy (anonymous):

yes i am sure

OpenStudy (anonymous):

can i go on or do u have more questions i dont mind

OpenStudy (anonymous):

No! Please, continue!

OpenStudy (anonymous):

okay sure will

OpenStudy (anonymous):

okay now this 1mole of O2= 32g 7*32=224 g 224 g of Oxygen reacts with 2 moles of ethane 62.3 g of oxygen =2*62.3/224=0.5 moles of ethane u got it

OpenStudy (anonymous):

Two questions. Where does 62.3 g of O2 come from?

OpenStudy (anonymous):

the 62.3 of O2 comes from just One molecule instead of 2

OpenStudy (anonymous):

AH! Yes, that makes sense. Carry on, love.

OpenStudy (anonymous):

what was ur other question

OpenStudy (anonymous):

Nevermind, it was lame :@

OpenStudy (anonymous):

hah okay

OpenStudy (anonymous):

So, is that it?

OpenStudy (anonymous):

gimmie a second

OpenStudy (anonymous):

:) As you wish.

OpenStudy (anonymous):

Do you have to explain it

OpenStudy (anonymous):

Unfortunately :/

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

which part do u want me to explain 1st

OpenStudy (anonymous):

@Brostep0s

OpenStudy (anonymous):

Actually, this will do fine. I think I can take it from here. Thanks. Hreat help, fit for gods. :*

OpenStudy (anonymous):

okayy if u need more help just tag me il be here

OpenStudy (anonymous):

Thanks again cx